OpenStudy (anonymous):

Solve 3x2 = -30x + 24. Using quadratic equation.

OpenStudy (anonymous):

first set it equal to 0 3x^2+30x-24=0 Then use the quadratic formula! a b and c are the coefficients so a=3, b=30 and c = -24 \[x=\frac{ -b \pm \sqrt{b^2-4ac} }{ 2a }\]

OpenStudy (anonymous):

\[x ^{2} + 10x -8 = 0\] \[\frac{ -10 \pm \sqrt{10^{2}+4*8} }{ 2 }\]

OpenStudy (anonymous):

So \[x = -5 \pm \sqrt{17}\] Have I done that right?

OpenStudy (anonymous):

I don't think so. I got \[-4\pm \sqrt{33}\] but thats not one of my answer choices..

OpenStudy (anonymous):

It should be -5 not -4. Check you work. http://www.wolframalpha.com/input/?i=3x2+%3D+-30x+%2B+24 Go to solutions.

OpenStudy (anonymous):

Yeah, wow I messed up my arithmetic there.